OpenStudy (across):

How do I solve x + 3 = 5? Now that that's out of the way, I quickly wanted to wish you guys a merry Christmas, a happy new year and a prosperous 2012. Enjoy your holidays to the fullest with family and friends. I've made great friendships here and think you guys are amazing. Keep up the good work, night owls!

OpenStudy (turingtest):

Merry ambiguous holidays from all!

OpenStudy (anonymous):

Merry Christmas and happy new year to you too.

OpenStudy (anonymous):

subtract 3 from each side x=2 HAppy Holidays to you, too!

OpenStudy (amistre64):

thanx, and same to you. whenever something is stuck by addition/subtraction I refer to it as a neighbor and we need to send the neighbor home. x+3 = 5 -3 -3 --------- x = 2 since the x has no "spouse" to get rid of we are done

OpenStudy (turingtest):

across goes to MIT, let's not insult her by answering her question

OpenStudy (across):

Haha! That's bound to leave a lasting impression, amistre! :)

OpenStudy (amistre64):

MITt romney?

OpenStudy (anonymous):

Btw excellent selection of the question lol :D

OpenStudy (mr.math):

Merry Christmas and happy new year to across and all!

OpenStudy (anonymous):

Same to you...

OpenStudy (anonymous):

and across, the answer to your question is x = 2.